# M3DHMR: Monocular 3D Hand Mesh Recovery

### v1

@@ Abstract
Monocular 3D hand mesh recovery is challenging due to high degrees of freedom of hands, 2D-to-3D ambiguity and self-occlusion. Most existing methods are either inefficient or less straightforward for predicting the position of 3D mesh vertices. Thus, we propose a new pipeline called Monocular 3D Hand Mesh Recovery (M3DHMR) to directly estimate the positions of hand mesh vertices. M3DHMR provides 2D cues for 3D tasks from a single image and uses a new spiral decoder consist of several Dynamic Spiral Convolution (DSC) Layers and a Region of Interest (ROI) Layer. On the one hand, DSC Layers adaptively adjust the weights based on the vertex positions and extract the vertex features in both spatial and channel dimensions. On the other hand, ROI Layer utilizes the physical information and refines mesh vertices in each predefined hand region separately. Extensive experiments on popular dataset FreiHAND demonstrate that M3DHMR significantly outperforms state-of-the-art real-time methods. The code will be available once the paper is accepted.
@@ 1 Introduction
Monocular 3D hand mesh recovery has attracted great attention in VR/AR, human-computer interaction, etc [30, 16]. This task aims to estimate the 3D spatial positions of hand mesh vertices from a single RGB image, which possesses broad commercial prospects with low-budget. However, high degrees of hand freedom, 2D to 3D ambiguity and self-occlusion make significant challenges for hand reconstruction [1, 9, 5].
To overcome these challenges, parametric model-based approaches [34, 6, 29, 27] usually use MANO [28] which obtains a hand mesh from coefficients of shape and pose to facilitately generate diverse hand poses for various applications with flexibility and high-precision. Though these methods are flexible, their 3D structure is not intuitive enough, especially the shape parameters are too difficult to optimize. Conversely, vertex-based methods can directly estimate the coordinates of the 3D mesh. Specifically, these models usually define the wrist of the hand as the root and calculate the root-relative 3D mesh coordinates. Some latest researches [22, 23, 18] are unable to implement in real-time due to their exclusive focus on improving the reconstruction accuracy, without considering the scale of models. Others [13, 20, 7, 33, 2, 30, 15, 31] can achieve real-time performance but are limited in reconstruction accuracy. Additionally, all methods previously mentioned have to provide the absolute root coordinate in camera space to get absolute camera coordinates of the 3D mesh, which limits the application to many high-level tasks. Though CMR [5] and MobRecon [3] provide pipelines to directly acquire the absolute 3D coordinates of the mesh in real-time, their reconstruction accuracy still needs to be improved.
@@ 2.1 Hand Mesh Estimation
Model-based approaches commonly rely on parametric hand models, such as MANO [28], which provides a low-dimensional representation of hand poses and shapes. Building upon this foundation, Zhou et al. [34] proposed the first real-time monocular hand capture framework, integrating multi-modal data (RGB, keypoints, and contours) to achieve highly accurate 3D reconstructions. Furthermore, Chen et al. [6] developed a self-supervised model-based 3D hand reconstruction method that eliminates the need for 3D annotation by leveraging geometric constraints and 2D-3D consistency for training. Meanwhile, RealisticHands [29] combined parametric mesh priors with non-parametric details, enabling high-quality hand pose and shape estimation from monocular RGB inputs. HaMeR [27] used a fully transformer-based architecture to analyze hands with significantly increased accuracy and robustness. Despite their efficiency, these methods are inherently limited by the expressiveness of parametric models.
Vertex-based methods directly predict the 3D coordinates of mesh vertices, offering flexibility beyond parametric models. Among these, Mesh Graphormer [23] and PointHMR [18] pioneered the use of transformer architectures for monocular 3D hand mesh recovery, achieving notable improvements in accuracy and robustness. However, their computational complexity hinders real-time deployment. To address this limitation, several efficient alternatives have emerged. MobileHand [13] introduced a lightweight network design, while YoutubeHand [20] and Pose2Mesh [7] leveraged graph convolutions for efficient mesh reconstruction. Further explorations include HIU-DMTL [33], which employed deep multi-task learning, and I2UV-HandNet [2], which adopted an image-to-UV mapping strategy. Tang et al. [30] enhanced real-time performance and significantly improved pose and shape accuracy through a novel alignment optimization module. CMR [5] and MobRecon [3] proposed end-to-end pipelines to directly estimate absolute 3D coordinates. More recently, CLIP-Hand3D [15] integrated vision-language pretraining with context-aware prompting to improve generalization in challenging scenarios, and FastViT [31] introduced a hybrid vision transformer with structural reparameterization for optimal speed-accuracy trade-offs.
@@ 2.2 Spiral Convolution
Defferrard et al. [8] pioneered graph convolutional networks (GCNs) that operate on non-Euclidean graph structures while maintaining computational complexity comparable to that of classical CNNs. Masci et al. [24] developed spatial graph convolution operators that incorporate geodesic-aware sampling of graph signals, providing more intuitive handling of 3D mesh vertices in the spatial domain. This line of research was further advanced by Lim et al. [21], who introduced spiral convolution based on graph convolution as an effective approach for processing 3D mesh vertices.
Subsequent improvements in mesh spatial convolution techniques include SpiralNet++ [10], which employed truncated spiral lines to limit sampled vertices while expanding the receptive field through hollow spiral lines. Kolotouros et al. [20] directly learned the mapping relationships from 2D images to 3D meshes by implementing a straightforward encoder-decoder architecture with spiral convolution. Recent works like CMR and MobRecon designed ISM and DSConv to further enhance feature extraction.
@@ 3.1 Overview
Taking a single RGB image as input, our goal is to get the 3D coordinates of the hand mesh vertices. M3DHMR firstly uses a two-stack hourglass network [26] to generate encoding features $F^{e}$ and predict 2D joint landmarks $L^{p}$. Note that the first stack and the first layer of the second stack are pretrained under guidance of heatmaps and landmarks. Moreover, M3DHMR adopts a spiral decoder to infer the root-relative coordinates of the 3D mesh vertices from the encoding features $F^{e}$. Specifically, pose features $F^{p}$ are first obtained by grid sampling:
@@
Subsequently, multiple graph-based DSC Layers produce multiple scale mesh vertex maps in a coarse-to-fine manner. Each layer upsamples output mesh features from the previous layer and explores spatial connectivity and channel dependencies between vertices to get fine-grained vertex features. Finally, M3DHMR adopts a ROI Layer to refine the result and uses the same method as CMR [5] to estimate the position of root in the camera space and thus recovers the coordinates of the 3D hand mesh vertices in the camera space. The overall network structure is shown in Fig. 2.
@@ 3.2 Dynamic Spiral Convolution Layer
Similar to SpiralConv [21] and SpiralConv++ [11], dynamic spiral convolution determines the convolution centre and produces a sequence of enumerated centre vertices based on adjacency, followed by the 1-ring vertices, the 2-ring vertices, and so on, until all vertices containing n rings are included:
@@ 3.3 Region of Interest Layer
The hand is artificially divided into six regions: palm, thumb, index finger, middle finger, ring finger, and little finger based on the shape and motion characteristics (See Fig. 3). We believe that the mesh vertices in the same region are highly correlated. The movement of the vertices strongly affects others within the same region but has less effect on vertices in other regions. For example, movement at the joint of the index finger causes significant displacement at the tip of the finger, but has minimal impact on the palm and other fingers. Therefore, we assume that for each vertex, the finger or palm region where it is located will be the region of interest. Benefiting from the large receptive field of dilated spiral convolution [11], ROI Layer can sufficiently explore the global features in each hand region of convolution centre to refine the mesh in predefined hand regions.
@@
Table 1: Ablation study of DSC Layer and ROI Layer with Kernel Size of 8.
3.4 Loss Function
We adopt $L_{1}$ loss to the 3D mesh loss $L_{mesh}$ and 2D pose loss $L_{pose2D}$. Inspired by [9], we adopt the normal loss $L_{norm}$ and the edge-length loss $L_{edge}$ to preserve the surface normals and penalize flying vertices. Additionally, we conduct consistency supervision in both 3D and 2D space similar to [14] and [32]. Formally, we have
$\mathcal{L}_{mesh}$ $=||V-V^{*}||_{1},$
$\mathcal{L}_{pose2D}$ $=||L^{p}-L^{p,*}||_{1},$
$\mathcal{L}_{norm}$ $=\sum_{c\in C}\sum_{(i,j)\subset c}|\frac{V_{i}-V_{j}}{||V_{i}-V_{j}||_{2}}\cdot n_{c}^{\star}|,$
${\mathcal{L}}_{edge}$ $=\sum_{c\in C}\sum_{(i,j)\subset c}|||V_{i}-V_{j}||_{2}-||V_{i}^{\star}-V_{j}^{\star}||_{2}|,$
$\mathcal{L}_{con3D}$ $=||V_{view1}-V_{view2}||_{1},$
$\mathcal{L}_{con2D}$ $=||L^{p}_{view1}-L^{p}_{view2}||_{1},$ (7)
@@ 3.4 Loss Function
SpiralConv++[5] 6.8 6.9 0.723 0.976
Depth-Separable SpiralConv[3] 6.8 6.9 0.724 0.977
Dynamic Spiral Convolution ($K$=8) 6.6 6.7 0.734 0.978
@@
4 Experiments
4.1 Datasets and Metric
We use the following metrics in quantitative evaluations.
PA-MPJPE/MPVPE ignores global variations and measures the mean per joint/vertex position error in terms of Euclidean distance (mm) between the root-relative prediction and the ground truth coordinates based on Procrustes analysis [12]. For brevity, the metric is abbreviated as PJ/PV.
FPS measures the number of frames an algorithm can process per second.
4.2 Experiment Setup
We set the batch size to 32 and train 38 epochs with Adam optimizer [19]. The learning rate is initialized to 0.001, which is divided by 10 at 30th epoch. Data augmentation includes random box scaling, box rotation, box shift and color jitter.
@@ 4.2 Experiment Setup
Non Real-Time Methods METRO [22] CVPR 21 HRNet 6.7 6.8 0.717 0.981 -
MeshGraphormer [23] ICCV 21 HRNet 5.9 6.0 0.764 0.986 -
PointHMR [18] CVPR 23 HRNet 6.1 6.6 0.720 0.984 -
HaMeR [27] CVPR 24 ViTPose 6.0 5.7 0.785 0.990 -
HaMeR-170k [27] CVPR 24 ViTPose 6.1 5.8 0.782 0.990 -
Real-Time Methods FreiHAND [35] ICCV 19 ResNet50 11.0 10.9 0.516 0.934 -
ObMan [17] CVPR 19 ResNet18 - 13.2 0.436 0.908 20
Boukhayma et al. [1] CVPR 19 ResNet50 - 13.0 0.435 0.898 -
MobileHand [13] ICONIP 20 MobileNetV3 - 13.1 0.439 0.902 -
YoutubeHand [20] CVPR 20 ResNet50 8.4 8.6 0.614 0.966 -
I2L-MeshNet [25] ECCV 20 ResNet50 7.4 7.6 0.681 0.973 33
Pose2Mesh [7] ECCV 20 HRNet 7.7 7.8 0.674 0.969 22
HIU-DMTL [33] ICCV 21 Hourglass 7.1 7.3 0.699 0.974 -
CMR [5] CVPR 21 ResNet50 6.9 7.0 0.715 0.977 30
I2UV-HandNet [2] ICCV 21 ResNet50 6.7 6.9 0.707 0.977 -
Tang et al. [30] ICCV 21 ResNet50 6.7 6.7 0.724 0.981 39
MobRecon [4] CVPR 22 GhostStack 8.8 9.1 0.597 0.960 -
MobRecon [4] CVPR 22 DenseStack 6.9 7.2 0.694 0.979 -
MobRecon [4] CVPR 22 ResNet18 6.7 6.8 0.727 0.979 69
MobRecon [4] CVPR 22 ResNet34 6.6 6.7 0.730 0.979 43
CLIP-Hand3D [15] ACM MM 23 ResNet50 6.6 6.7 0.728 0.981 77
FastViT [31] ICCV 23 FastViT 6.6 6.7 0.722 0.981 -

### v2

@@ Abstract
Monocular 3D hand mesh recovery is challenging due to high degrees of freedom of hands, 2D-to-3D ambiguity and self-occlusion. Most existing methods are either inefficient or less straightforward for predicting the position of 3D mesh vertices. Thus, we propose a new pipeline called Monocular 3D Hand Mesh Recovery (M3DHMR) to directly estimate the positions of hand mesh vertices. M3DHMR provides 2D cues for 3D tasks from a single image and uses a new spiral decoder consist of several Dynamic Spiral Convolution (DSC) Layers and a Region of Interest (ROI) Layer. On the one hand, DSC Layers adaptively adjust the weights based on the vertex positions and extract the vertex features in both spatial and channel dimensions. On the other hand, ROI Layer utilizes the physical information and refines mesh vertices in each predefined hand region separately. Extensive experiments on popular dataset FreiHAND demonstrate that M3DHMR significantly outperforms state-of-the-art real-time methods. The code is available at https://github.com/Jackson-coder/M3DHMR.
@@ 1 Introduction
Monocular 3D hand mesh recovery has attracted great attention in VR/AR, human-computer interaction, etc [tang2021towards, 2020MEgATrack]. This task aims to estimate the 3D spatial positions of hand mesh vertices from a single RGB image. It improves the accuracy and security of hand gesture authentication by restoring biometric features. On the one hand, it repairs low-quality samples caused by factors such as wear, contamination, or limitations of capture devices, which reduces false recognition rates. On the other hand, it reconstructs incomplete data, thereby boosting performance of multimodal hand gesture authentication system. However, high degrees of hand freedom, 2D to 3D ambiguity and self-occlusion make significant challenges for hand reconstruction [boukhayma20193d, ge20193d, Chen_Liu_Ma_Chang_Wang_Chen_Guo_Wan_Zheng_2021].
To overcome these challenges, parametric model-based approaches [zhou2020monocular, chen2021model, seeber2021realistichands, pavlakos2024reconstructing] usually use MANO [Romero_Tzionas_Black_2017] which obtains a hand mesh from coefficients of shape and pose to facilitately generate diverse hand poses for various applications with flexibility and high-precision. Though these methods are flexible, their 3D structure is not intuitive enough, especially the shape parameters are too difficult to optimize. Conversely, vertex-based methods can directly estimate the coordinates of the 3D mesh. Specifically, these models usually define the wrist of the hand as the root and calculate the root-relative 3D mesh coordinates. Some latest researches [lin2021end, lin2021-mesh-graphormer, PointHMR] are unable to implement in real-time due to their exclusive focus on improving the reconstruction accuracy, without considering the scale of models. Others [MobileHand:2020, kulon2020weakly, Choi_2020_ECCV_Pose2Mesh, zhang2021hand, chen2021i2uv, tang2021towards, guo_clip_hand, vasufastvit2023] can achieve real-time performance but are limited in reconstruction accuracy. Additionally, all methods previously mentioned have to provide the absolute root coordinate in camera space to get absolute camera coordinates of the 3D mesh, which limits the application to many high-level tasks. Though CMR [Chen_Liu_Ma_Chang_Wang_Chen_Guo_Wan_Zheng_2021] and MobRecon [Chen_Liu_Dong_Zhang_Ma_Xiong_Zhang_Guo_2022] provide pipelines to directly acquire the absolute 3D coordinates of the mesh in real-time, their reconstruction accuracy still needs to be improved.
@@ 2.1 Hand Mesh Estimation
Model-based approaches commonly rely on parametric hand models, such as MANO [Romero_Tzionas_Black_2017], which provides a low-dimensional representation of hand poses and shapes. Building upon this foundation, Zhou et al. [zhou2020monocular] proposed the first real-time monocular hand capture framework, integrating multi-modal data (RGB, keypoints, and contours) to achieve highly accurate 3D reconstructions. Furthermore, Chen et al. [chen2021model] developed a self-supervised model-based 3D hand reconstruction method that eliminates the need for 3D annotation by leveraging geometric constraints and 2D-3D consistency for training. Meanwhile, RealisticHands [seeber2021realistichands] combined parametric mesh priors with non-parametric details, enabling high-quality hand pose and shape estimation from monocular RGB inputs. HaMeR [pavlakos2024reconstructing] used a fully transformer-based architecture to analyze hands with significantly increased accuracy and robustness. Despite their efficiency, these methods are inherently limited by the expressiveness of parametric models.
Vertex-based methods directly predict the 3D coordinates of mesh vertices, offering flexibility beyond parametric models. Among these, Mesh Graphormer [lin2021-mesh-graphormer] and PointHMR [PointHMR] pioneered the use of transformer architectures for monocular 3D hand mesh recovery, achieving notable improvements in accuracy and robustness. However, their computational complexity hinders real-time deployment. To address this limitation, several efficient alternatives have emerged. MobileHand [MobileHand:2020] introduced a lightweight network design, while YoutubeHand [kulon2020weakly] and Pose2Mesh [Choi_2020_ECCV_Pose2Mesh] leveraged graph convolutions for efficient mesh reconstruction. Further explorations include HIU-DMTL [zhang2021hand], which employed deep multi-task learning, and I2UV-HandNet [chen2021i2uv], which adopted an image-to-UV mapping strategy. Tang et al. [tang2021towards] enhanced real-time performance and significantly improved pose and shape accuracy through a novel alignment optimization module. CMR [Chen_Liu_Ma_Chang_Wang_Chen_Guo_Wan_Zheng_2021] and MobRecon [Chen_Liu_Dong_Zhang_Ma_Xiong_Zhang_Guo_2022] proposed end-to-end pipelines to directly estimate absolute 3D coordinates. More recently, CLIP-Hand3D [guo_clip_hand] integrated vision-language pretraining with context-aware prompting to improve generalization in challenging scenarios, and FastViT [vasufastvit2023] introduced a hybrid vision transformer with structural reparameterization for optimal speed-accuracy trade-offs.
@@ 2.2 Spiral Convolution
Defferrard et al. [defferrard2016convolutional] pioneered graph convolutional networks (GCNs) that operate on non-Euclidean graph structures while maintaining computational complexity comparable to that of classical CNNs. Masci et al. [masci2015geodesic] developed spatial graph convolution operators that incorporate geodesic-aware sampling of graph signals, providing more intuitive handling of 3D mesh vertices in the spatial domain. This line of research was further advanced by Lim et al. [lim2018simple], who introduced spiral convolution based on graph convolution as an effective approach for processing 3D mesh vertices.
Subsequent improvements in mesh spatial convolution techniques include SpiralNet++ [gong2019spiralnet++], which employed truncated spiral lines to limit sampled vertices while expanding the receptive field through hollow spiral lines. Kolotouros et al. [kulon2020weakly] directly learned the mapping relationships from 2D images to 3D meshes by implementing a straightforward encoder-decoder architecture with spiral convolution. Recent works like CMR and MobRecon designed ISM and DSConv to further enhance feature extraction.
@@ 3.1 Overview
Taking a single RGB image as input, our goal is to get the 3D coordinates of the hand mesh vertices. M3DHMR firstly uses a two-stack hourglass network [newell2016stacked] to generate encoding features $F^{e}$ and predict 2D joint landmarks $L^{p}$. Note that the first stack and the first layer of the second stack are pretrained under guidance of heatmaps and landmarks. Moreover, M3DHMR adopts a spiral decoder to infer the root-relative coordinates of the 3D mesh vertices from the encoding features $F^{e}$. Specifically, pose features $F^{p}$ are first obtained by grid sampling:
@@
Subsequently, multiple graph-based DSC Layers produce multiple scale mesh vertex maps in a coarse-to-fine manner. Each layer upsamples output mesh features from the previous layer and explores spatial connectivity and channel dependencies between vertices to get fine-grained vertex features. Finally, M3DHMR adopts a ROI Layer to refine the result and uses the same method as CMR [Chen_Liu_Ma_Chang_Wang_Chen_Guo_Wan_Zheng_2021] to estimate the position of root in the camera space and thus recovers the coordinates of the 3D hand mesh vertices in the camera space. The overall network structure is shown in Fig. 2.
@@ 3.2 Dynamic Spiral Convolution Layer
Similar to SpiralConv [lim2018simple] and SpiralConv++ [Gong_Chen_Bronstein_Zafeiriou_2019], dynamic spiral convolution determines the convolution centre and produces a sequence of enumerated centre vertices based on adjacency, followed by the 1-ring vertices, the 2-ring vertices, and so on, until all vertices containing n rings are included:
@@ 3.3 Region of Interest Layer
The hand is artificially divided into six regions: palm, thumb, index finger, middle finger, ring finger, and little finger based on the shape and motion characteristics (See Fig. 3). We believe that the mesh vertices in the same region are highly correlated. The movement of the vertices strongly affects others within the same region but has less effect on vertices in other regions. For example, movement at the joint of the index finger causes significant displacement at the tip of the finger, but has minimal impact on the palm and other fingers. Therefore, we assume that for each vertex, the finger or palm region where it is located will be the region of interest. Benefiting from the large receptive field of dilated spiral convolution [Gong_Chen_Bronstein_Zafeiriou_2019], ROI Layer can sufficiently explore the global features in each hand region of convolution centre to refine the mesh in predefined hand regions.
3.4 Loss Function
We adopt $L_{1}$ loss to the 3D mesh loss $L_{mesh}$ and 2D pose loss $L_{pose2D}$. Inspired by [ge20193d], we adopt the normal loss $L_{norm}$ and the edge-length loss $L_{edge}$ to preserve the surface normals and penalize flying vertices. Additionally, we conduct consistency supervision in both 3D and 2D space similar to [guo2018stacked] and [yang2021semihand]. Formally, we have
[…]
4 Experiments
4.1 Datasets and Metric
We use the following metrics in quantitative evaluations.
PA-MPJPE/MPVPE ignores global variations and measures the mean per joint/vertex position error in terms of Euclidean distance (mm) between the root-relative prediction and the ground truth coordinates based on Procrustes analysis [gower1975generalized]. For brevity, the metric is abbreviated as PJ/PV.
FPS measures the number of frames an algorithm can process per second.
4.2 Experiment Setup
We set the batch size to 32 and train 38 epochs with Adam optimizer [kingma2014adam]. The learning rate is initialized to 0.001, which is divided by 10 at 30th epoch. Data augmentation includes random box scaling, box rotation, box shift and color jitter. We use PyTorch as the framework to conduct all experiments and use a single NVIDIA RTX 3090 for training and a single NVIDIA RTX 2080 Ti for inference.
@@ 4.2 Experiment Setup
Table 1: Ablation study of DSC Layer and ROI Layer with Kernel Size of 8. ”Baseline” denotes the model with only Spiral Convolution Layers.
@@
SpiralConv++[Chen_Liu_Ma_Chang_Wang_Chen_Guo_Wan_Zheng_2021] 6.8 6.9 0.723 0.976
Depth-Separable SpiralConv[Chen_Liu_Dong_Zhang_Ma_Xiong_Zhang_Guo_2022] 6.8 6.9 0.724 0.977
Dynamic Spiral Convolution (KK=8) 6.6 6.7 0.734 0.978
@@
Non Real-Time
Methods
METRO [lin2021end] CVPR 21 HRNet 6.7 6.8 0.717 0.981 -
MeshGraphormer [lin2021-mesh-graphormer] ICCV 21 HRNet 5.9 6.0 0.764 0.986 -
PointHMR [PointHMR] CVPR 23 HRNet 6.1 6.6 0.720 0.984 -
HaMeR [pavlakos2024reconstructing] CVPR 24 ViTPose 6.0 5.7 0.785 0.990 -
HaMeR-170k [pavlakos2024reconstructing] CVPR 24 ViTPose 6.1 5.8 0.782 0.990 -
Real-Time
Methods
FreiHAND [zimmermann2019freihand] ICCV 19 ResNet50 11.0 10.9 0.516 0.934 -
ObMan [hasson19_obman] CVPR 19 ResNet18 - 13.2 0.436 0.908 20
Boukhayma et al. [boukhayma20193d] CVPR 19 ResNet50 - 13.0 0.435 0.898 -
MobileHand [MobileHand:2020] ICONIP 20 MobileNetV3 - 13.1 0.439 0.902 -
YoutubeHand [kulon2020weakly] CVPR 20 ResNet50 8.4 8.6 0.614 0.966 -
I2L-MeshNet [moon2020i2l] ECCV 20 ResNet50 7.4 7.6 0.681 0.973 33
Pose2Mesh [Choi_2020_ECCV_Pose2Mesh] ECCV 20 HRNet 7.7 7.8 0.674 0.969 22
HIU-DMTL [zhang2021hand] ICCV 21 Hourglass 7.1 7.3 0.699 0.974 -
CMR [Chen_Liu_Ma_Chang_Wang_Chen_Guo_Wan_Zheng_2021] CVPR 21 ResNet50 6.9 7.0 0.715 0.977 30
I2UV-HandNet [chen2021i2uv] ICCV 21 ResNet50 6.7 6.9 0.707 0.977 -
Tang et al. [tang2021towards] ICCV 21 ResNet50 6.7 6.7 0.724 0.981 39
MobRecon [chen2022mobrecon] CVPR 22 GhostStack 8.8 9.1 0.597 0.960 -
MobRecon [chen2022mobrecon] CVPR 22 DenseStack 6.9 7.2 0.694 0.979 -
MobRecon [chen2022mobrecon] CVPR 22 ResNet18 6.7 6.8 0.727 0.979 69
MobRecon [chen2022mobrecon] CVPR 22 ResNet34 6.6 6.7 0.730 0.979 43
CLIP-Hand3D [guo_clip_hand] ACM MM 23 ResNet50 6.6 6.7 0.728 0.981 77
FastViT [vasufastvit2023] ICCV 23 FastViT 6.6 6.7 0.722 0.981 -